OpenStudy (anonymous):

a soultion of ammonia and water contains 4.00 * 10^25 water mol and 6.80*10^24 ammonia mol . How many hydrogen atoms are in this solution?

OpenStudy (anonymous):

We need to find the amount of hydrogen atoms in water and in ammonia. we'll do them one at a time and add the results after. I'lll do water and I'd like to see you attempt Ammonia \[Water: H_20\]\[Ammonia: NH_3\] the percentage of hydrogen atoms in the water molecule is equal to the percentage of mass that hydrogen has vs the entire molecule's mass. we will use their molar masses from the periodic table: \[FractionHydrogen = \frac{ Mass Hydrogen }{ Mass H_20 } = \frac{ 2 \times 1.008 }{ (2 \times 1.008) + 16 } = 0.112\] So 11.2% of 4.00*10^25 mol are hydrogen. \[0.112 \times 4\times10^{25} = 4.48 \times 10^{24} mol Hydrogen\] use avogadro's number to find how many atoms in that many mols of hydrogen. I get 2.7*10^48 I have to go eat but I'll check soon

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."